“Charley Chicken” and “Bradley Bee” are brands of canned tuna. During a week a certain amount of advertising appears for these products. There may be no advertising, one form of advertising (newspaper coupon), or two forms of advertising (coupon and a special store display). Let C denote the level of advertising for Charley Chicken. It can take the values c=0, 1 or 2. Let B denote the level of advertising for Bradley Bee; B can take the values b=0, 1 or 2. Suppose the following table represents the joint probability distribution of the advertising levels for these two brands of canned tuna.

Bradley Bee’s Advertising: B

0

1

2

0

0.05

0.15

0.25

Charley Chicken’s Advertising: C

1

0.05

0.05

0.05

2

0.10

0.15

0.15

Find the marginal probability f(b=1).

Options:

F(b=1)=0.25

F(b=1)=0.30

F(b=1)=0.35

F(b=1)=0.40

F(b=1)=0.45
Answer #1

Solution:

The marginal probability is calculated as follows:

F(b=1)= P(b =1,c=0) + P(b=1,c=1) + P(b=1,c=2)

= 0.15 + 0.05 + 0.15

= 0.35

So option 3 is the correct answer.